# RSCC: A Large-Scale Remote Sensing Change Caption Dataset for Disaster Events

### v2

@@ 6 Limitations
Due to the lack of proficient labels and complexity of image pairs themselves, the generated captions may contains vague descriptions which is even hard for experts to clarify. Besides, we only employ text similarity metrics because existing image–to-text captioning metrics (e.g., FLEUR (Lee et al., ), SPARC (Jung et al., ) and G-VEval (Tong et al., )) only focus on single image which fail in multi-image scenario. We leave these parts for future work.
@@ A.1 RSCC Captioning Details
We compare the performance of large-size MLLMs with zero-shot template (A.6) including open-source models such as Pixtral Large (Mistral AI Team, ) and LLaMA-4 Maverick (Meta AI, ). We also conduct case study on proprietary models including GPT-4.1 (2025-04-14) (OpenAI, ), Gemeni-2.5-Pro (2025-03-25) (Gemini Team, ), and Qwen-VL Max (2025-01-25) (Qwen Team, ), along with reasoning model such as Claude-3.7-Sonnet-Thinking (2025-02-25) (Anthropic AI, ) and o3 (2025-04-03)  (OpenAI, ). For results generation, We use default configurations of the above models. Figure 89 and 10 show qualitative results of empirical study. We found proprietary models outperform open-sourced models in completeness and accuracy. Visual reasoning notably improve the quality of caption in completeness but it also introduce vague information even hallucinations. As remote sensing change captioning requires world knowledge and complex reasoning, the latest state-of-the-art MLLMs seem to be insufficient.
@@ A.2 More Results
In Scenario (c), ground truth provides the most accurate, complete, and factually consistent description. It captures the catastrophic scale of destruction ("every structure reduced to rubble," "barren landscape"), explicitly mentions debris and empty streets, and aligns with typical patterns of severe wind-driven disasters (e.g., hurricanes or tornadoes). While it does not specify the disaster type, its focus on observable damage patterns (total structural collapse, vegetation loss) adheres strictly to visual evidence. Other captions either misinterpret the scene (Phi-4-$M$, CCExpert), lack detail (TEOChat), or omit critical damage indicators (Qwen2-VL).
@@
Figure 12 shows more samples on RSCC subset along with baseline results. Table 3 shows overall quantitative results on RSCC subset. It is witnessed that auxiliary building damage info augmentation greatly improve the quality of change captions. We also find performance gets saturated equipped with auxiliary info regardless model size. We provide an additional metric BLEURT  [^4] https://huggingface.co/lucadiliello/BLEURT-20-D12 (Sellam et al., 2020), a learned evaluation metric to measure contextual similarity as well. However, the BLEURT is strongly biased on text length, which fails in valid evaluation. We are seeking for more reliable metrics in the future. Table 4 and  5 display RSCC data source details and baseline model configurations respectively.
@@
LLaVA-NeXT-Interleave (8B) (Li et al., 2024b) 12.51 13.29 42.80 46.99 57
@@ A.4 Author Statement and Data License
Data License Confirmation: The dataset is released under the [$C$-BY-4.0], which permits unrestricted use, distribution, and reproduction in any medium, provided the original work is properly cited.

### v4

@@ 6 Limitations
Due to the lack of proficient labels and complexity of image pairs themselves, the generated captions may contains vague descriptions which is even hard for experts to clarify. Besides, we only employ text similarity metrics because existing image–to-text captioning metrics (e.g., FLEUR [28], SPARC [26] and G-VEval [67]) only focus on single image which fail in multi-image scenario. We leave these parts for future work.
@@ A.1 RSCC Captioning Details
We compare the performance of large-size MLLMs with zero-shot template (A.6) including open-source models such as Pixtral Large [47] and LLaMA-4 Maverick [44]. We also conduct case study on proprietary models including GPT-4.1 (2025-04-14) [50], Gemeni-2.5-Pro (2025-03-25) [16], and Qwen-VL Max (2025-01-25) [55], along with reasoning model such as Claude-3.7-Sonnet-Thinking (2025-02-25) [1] and o3 (2025-04-03)  [49]. For results generation, We use default configurations of the above models. Figure 89 and 10 show qualitative results of empirical study. We found proprietary models outperform open-sourced models in completeness and accuracy. Visual reasoning notably improve the quality of caption in completeness but it also introduce vague information even hallucinations. As remote sensing change captioning requires world knowledge and complex reasoning, the latest state-of-the-art MLLMs seem to be insufficient.
@@ A.2 More Results
In Scenario (c), ground truth provides the most accurate, complete, and factually consistent description. It captures the catastrophic scale of destruction ("every structure reduced to rubble," "barren landscape"), explicitly mentions debris and empty streets, and aligns with typical patterns of severe wind-driven disasters (e.g., hurricanes or tornadoes). While it does not specify the disaster type, its focus on observable damage patterns (total structural collapse, vegetation loss) adheres strictly to visual evidence. Other captions either misinterpret the scene (Phi-4-MM, CCExpert), lack detail (TEOChat), or omit critical damage indicators (Qwen2-VL).
@@
Figure 12 shows more samples on RSCC subset along with baseline results. Table 3 shows overall quantitative results on RSCC subset. It is witnessed that auxiliary building damage info augmentation greatly improve the quality of change captions. We also find performance gets saturated equipped with auxiliary info regardless model size. We provide an additional metric BLEURT  [^4] https://huggingface.co/lucadiliello/BLEURT-20-D12 [62], a learned evaluation metric to measure contextual similarity as well. However, the BLEURT is strongly biased on text length, which fails in valid evaluation. We are seeking for more reliable metrics in the future. Table 4 and  5 display RSCC data source details and baseline model configurations respectively.
@@
LLaVA-NeXT-Interleave (8B) [31] 12.51 13.29 42.80 46.99 57
@@ A.4 Author Statement and Data License
Data License Confirmation: The dataset is released under the [CC-BY-4.0], which permits unrestricted use, distribution, and reproduction in any medium, provided the original work is properly cited.